OpenStudy (anonymous):

If you are finding the line perpendicular to x + 3y = 5, what is the slope of the perpendicular line? A. -3/5 B. 1/3 C. 3 D. -3

OpenStudy (anonymous):

If you are finding the line perpendicular to x + 3y = 5, what is the slope of the perpendicular line? A. -3/5 B. 1/3 C. 3 D. -3 please help

OpenStudy (anonymous):

are you there?

OpenStudy (anonymous):

hello?

OpenStudy (piasolywoda):

When two lines are perpendicular, the product of their slopes is -1. To answer this question, find the slope of the linear equation given; To do this, solve it for y; 3y = -x + 5; divide both sides by 3: y = -1/3 x + 5/3; and the slope is the coefficient of the variable x. In this case, the slope of the given line is -1/3. The slope of the line perpendicular to the given line is found by computing the negative reciprocal of -1/3. Look at your list of options and decide which one is the negative reciprocal of -1/3 and you should conclude that C is the correct answer.

OpenStudy (anonymous):

thank you. that really explains it